OpenStudy (anonymous):

write a system of equations with the solution (4, -3)

OpenStudy (texaschic101):

your equations are : x + y = 1 and x - y = 7 x + y = 1 --> x = 1 - y now we have : x - y = 7 (1 - y) - y = 7 -2y + 1 = 7 -2y = 7 - 1 -2y = 6 y = -3 now sub -3 in for y in either of the equations x + y = 1 x - 3 = 1 x = 3 + 1 x = 4 there you go....solved by substitution